Metop-2/A, All Instruments, Instrument Turn On Update, 24 April 2007 - Issued: 24 April 2007; 1140 UTC

Metop Instruments Turn On update

Date/Time(UTC) Message Issued:
24 April 2007, 1140 UTC

Satellite(s) Involved:

Instrument(s) Involved:
All instruments

Product(s) Involved/Affected:
All Products

Date/Time(UTC) of Initial Implementation:
24 April 2007, Time 0837 UTC

Details/Specifics of Change:
Eumetsat has issued the following notice of planned activities for week 17/2007 of Metop-2/A commissioning activities:
Commissioning activities continue and are due to complete during May 2007.
Commissioning activities will present a higher risk of limited service interruptions during week 17.
Following the total switch off of the MetOp-A payload module including all instruments on 20/4/2007, the investigation to understand the root cause is proceeding. In the mean time, the operational recovery started on the 23 April.

The following are the indicative dates for the switch-on of the instruments and HRPT. These dates could however be subject to slight variation depending on the achieved thermal conditions on board the MetOp-A satellite.

Instrument switch-on dates (scheduled):
AVHRR: Back in measurement mode on 24/4 (latest 25/4 a.m. for IR channels)
HIRS: Back in measurement mode on 24/4 (latest 25/4 a.m)
MHS: Back to measurement mode on 24/4 - Data nominal after thermal stabilisation (latest 25/4)
AMSU: Back to measurement mode on 24/4 - followed by two days of thermal stabilisation - Nominal status expected around 26/4
IASI: Back to measurement mode on 30/4
ASCAT: Back to measurement mode on 25/4
GOME: Back to measurement mode and standard timeline activation on 26/4
GRAS: Back to measurement mode (occultation) on 25/4
SEM: Back to measurement mode on 24/4
A-DCS: Back to measurement mode with operational configuration 25/4
S&R: Back to operations on 24/4
HRPT: Switch on 24/4 (sending filler data initially, until instruments are switched on)

Dissemination Restart Dates:
After an instrument is returned to measurement mode, a few orbits must be used to ensure that the product quality is still nominal. Dissemination of the Level 1 data can then be resumed. For the instruments requiring thermal stabilisation, the start of dissemination may require 1-2 additional days.

Specific announcements will be made to the operational users in advance of the resumption of the data and products dissemination.
Contact Person(s) Name/Email/Phone Number for Questions:
Emily Harrod 301-817-3882
Renee Smith Dearing 301-817-4117

Web site(s) that contain other relevant information (where applicable): - History of level 1B notices - All user notices - SOCC polar satellite status - SOCC geostationary satellite status - SOCC morning reports

* Emily D. Harrod
* Computer Operations Branch
* Pre-Product Processing Group
* E/SP11, NSOF, room 1460
* 4231 Suitland Road
* Suitland, Maryland 20746-4304
* voice 301-817-3882
* fax 301-817-3902
* email